OpenStudy (anonymous):

Solve: −negative 1 over 2 x + 3 = −x + 7 A) 1/2 B) 8 C) -8 D) 2

OpenStudy (anonymous):

\[\frac{-1}{2x+3}=-x+7\] First thing to do is to bring the \(2x+3\) up from the bottom of the fraction. What do you get when you multiply both sides by \(2x+3\)?

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

Nope, when you multiply both sides by \(2x+3\) you get: \[-1=(2x+3)(-x+7)\] Can you expand the brackets on the right hand side?

OpenStudy (anonymous):

no you cant?

OpenStudy (anonymous):

Yes: \[-1=-2x^2-3x+14x+21\] Take everything over to the left hand side: \[2x^2-11x-20=0\] Now you have a quadratic equation in standard form. Are you able to solve this now?

OpenStudy (anonymous):

so the answer is 8